Calculus I, Fall 2011 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 09

Assigned Nov 10, 2011.

1. Section 5.6: Problems 53, 55, 56, 58, 67, 71, 73.

2. Section 6.1: Problems 45, 47, 59, 65.

3. Section 6.2: Problems 28(c), 32, 45, 46.

4. Section 6.3: Problems 41, 50, 51, 57, 65, 77, 87, 89, 91.

5. Section 6.3: Do as many as time permits in problems 31-40, 85-90.